OpenStudy (anonymous):

Solve for x. 15 ≥ -3x or 2/5 x ≥ -2 {x | x ≤ -5} {x | x ≥ -5} {all reals}

OpenStudy (anonymous):

so first lets solve 15>= - 3x divide both sides by -3 and redo the sign >= into =< -5=<x now about the second 2/5x >= -2 multiply eah side by 5/2 x>=-2*5/2 x>=-5

OpenStudy (anonymous):

both equations gave the same answer, so the final answer will be x bigger or equal to -5.

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

np :)